DEIMOS shipping news - historical

Dave's report, April 17, 2002: We installed the new renishaw encoder mount on the rotation drive. DEIMOS is now rotating under computer control. It is not yet controlled by DCS. One of the cable wrap tensoners failed. It will need to be replaced when we find the spare. Jeff and Matt also started installing the forward cladding. The new unload positions for the gratings were found and the information was given to Dean to update the code. Barry worked on finding the problem with the thermal cutout for the AC to the carriage. Bob and Chris did further work to determine the source of the fixed pattern noise in the FCS CCD readout. The empty container arrived, which we will use to take back Lick commissioning materials that are no longer needed at Keck. Will arrived today.

Dave's report, April 16, 2002: We adjusted the patch of the drive disk. It is as good as we have seen. The skew of the drive rollers will be worked on later. We fixed the tilt drive of slider 4 and installed the 600 line grating in slider 4. We installed all the science filters in their correct positions. The camera was shimmed an additional 0.020 inches. Troubleshooting of the rotational drive system was begun. Chris Wright and Bob Kibrick spent time investigating the fixed pattern noise in the FCS images.

Dave's report, April 15, 2002: We installed the rotation drive servo motor and adjusted the balance. Barry worked on wiring issues in the carriage. At the end of the day, DEIMOS was rotating under power, in manual control. The 0.020 shims in the camera were installed. Work was done on the grating transport mechanism limits, and we loaded the 900 line grating into slider 3. We still have problems going to the unload position for slider 3. During the day we had some problems with the dewar chamber ion pump tripping out, probably with something out-gassing. Last night we saw the first spectra taken with DEIMOS on the summit.

Dave's report April 12, 2002: The chamber pumped down to 4 x 10-6 Torr. Unfortunately when removing it from the pump the "plug" in the dewar was not properly seated and the chamber vented to atmosphere. Later in the day we pumped the chamber and got it down to 3.6 x 10-5 Torr when we elected to pump the junction and the cold can and cool the dewar assembly. We started cooling at 5:00. On Saturday the chamber got down to -129C which is colder than we have ever seen the dewar. The tent mirror was installed. The piezo was activated and released to De Clark. The dewar stages were connected to the Galil, checked out, and released to De. Various temperature reading difficulties were tracked down and corrected, but some still remain.

Dave's report, April 9th, 2002: We had some continuing problems with the filter wheel, but they were corrected. It was released to De Clarke for testing. Dick Kanto installed the filter wheel cover. Preparations were made to mount the tent mirror and the dummy. The CCD controllers were mounted in the instrument along with the UPS, and communications was established to them. Bob and Chris completed their tests of the dewar system, and it was allowed to warm up. We expect it to be at room temperature by tomorrow morning.

Dave's report, April 8th, 2002: The filter wheel and shutter were installed. A fuducial on the filter wheel was damaged and was replaced. The thermal sensors on the camera were also installed. Jim Ward installed the FC fiber optics. Chris and Bob completed the tests they needed to do and by the end of the day they were prepared to let the dewar start to warm up. Preparations to install the dewar on Thursday are in progress. The biggest problem so far is that we do not have all the screws to hold the field flatner on to the dewar. The screws which were in Santa Cruz are being Fed Exed to arrive in Hawaii Wednesday. Clamp 5 on the grating system does not always indicate it is engaged and needs to be checked. This condition occurs only when selecting slider 3.

Dave's report, April 5th, 2002: The camera was installed in the spectrograph! The tilt will be aligned as well as we can do it with the instrument in the storage position. This may have to be redone when we are on the kinematics. Bob, Chris and Terry unpacked the controllers and VME crates and started wiring them up and checking them out. By the end of the day Chris has tested all the signals and Bob had a warm image. The dewar was filled with LN2 to cool overnight. Grant Tollith got the first image from the TV camera.

Dave's report, April 4th, 2002: The last air shipment arrived which was the dewar controllers and VME crates. Bob Kibrick and Chris Wright arrived with the shipment. The shipment was taken to the summit and stored in the instrument lab. Ken and De checked various communication elements between the software and the hardware. DEIMOS was moved into the observing position to check the operation of the switches that sense where on the telescope the instrument is. Some changes need to be made to the logic or wiring to the switches. The instrument clears the telescope by about inches when in the observing position. Vern completed the design of the renishaw encoder mounts that we agreed to change. The new parts have been fabricated and will arrive on Monday with Jim Ward.

Dave's report, April 3, 2002: We installed the collimator in DEIMOS! The camera cradle was aligned to the numbers it was measured to be at in Santa Cruz. We torqued the bolts holding the cradle to 300 ft-lbs. Assembly of the detector chamber to the cold can was completed, and the initial testing of the system is being done in the instrument lab. Today, the dewar system was bolted to the white cart, and Terry connected the cold can to the detector housing. Chamber and connection were pumped.

Dave's report, March 26-29th, 2002: On Wednesday the 24th the computers arrived with De Clarke. They were physically installed on Thursday, and on-line Friday. Currently, Celeste, Deitrader, and Roto, all with new Hawaiian names, are on-line and communicating with the instrument and the world. Barry went home on Wednesday. Ken tuned the flow meter to the Photometrics camera, and it now correctly reads flow to the camera head. Ken also connected the laptop computer to help with testing of the slitmask system. Slit masks were successfully installed from all slots. However, occasionally a mask would foul the pickoff mirror cell. This problem is likely the guide near the pickoff mirror cell that is intended to deflect the mask around the cell. Masks that are following the slitmask frame very closely can get between the guide and the frame and then foul the cell. A fix was tried that worked for many insertions, but then failed. This fix will be improved this week.

Dave's report, March 22, 2002: Jeff, Matt and Dick have the grating transport drive system operating and mechanically tested. The sliders clamp successfully (manually), and the fiducials are installed and adjusted. The pneumatic drive for the front hatch operates, as does the slitmask air cylinders. They mounted the science filter wheel and had it running using the control system. The filter wheel was subsequently removed for installation of the camera. Barry, Ken and Carol have tested the grating and slitmask systems using low level control, and they operate correctly. They have the new numbers for the grating clamping. They have also run the TV stages and the science filter wheel. The calibration lamps are installed and tested also. Sandy, Vern and David H completed testing of the camera and with Jeff's crew, and it was put in the storage box to await installation. The results of the camera testing are very encouraging, They are shown separately on Drew Phillip's web site.

Dave's report, March 18, 2002: Barry, Carol and Ken are moving stages under low level control. The TV stages have been checked and work properly. A motor that represents the servo for the dewar focus and translation stages, along with one that represents the filter wheel servo, were all driven with low level code. The slitmask scissor jack was driven and sees its limits. However, there are some open wires on the slitmask status bits that are a bit of a puzzle.

Jeff and crew helped mount the camera in the test stand in the dome for cold testing. They also continued on with a number of small projects, including preparing to install the baffling. Compressed air was connected to the instrument, and they are prepared to start testing the solenoids as soon as they have time on the instrument. Vernon and David Hilyard started cold testing the camera and did motion tests today. Tomorrow, when Sandy is available, they plan to do the image quality tests.

Dave's report, March 15, 2002: Electronics has now connected all the cables to the nose area mechanisms and is preparing to test them using low level code and the laptops. The Instrument group did a number of small things and are also prepared for the testing of mechanisms starting on Monday. They helped remove the camera from the test stand and put it in its box. The camera testing continued in the instrument lab. Analysis done by Drew Phillips of the images show that the image quality is significantly better than seen in Santa Cruz. This is the expected and hoped for result of the camera being cooler than it was in Santa Cruz. The camera was designed to work at 0C. The temperature as tested in Santa Cruz was about 21C and as tested in the instrument lab on the summit, 13C. The camera was removed from the test stand and put in its athermalizing box at the end of the day. Camera and box are in the dome for the weekend. On Monday we plan to mount the camera in the test stand and continue testing there.

Dave's report, March 14: The camera testing has started today in the instrument lab. Initial results of the image motion tests show that it is the same as measured in Santa Cruz (6 pixels). Jim Ward assembled Joe's telescope, and they are using it in the instrument lab to test image quality. Cold testing is expected to start on Monday after the camera has athermalized over the weekend. Jeff and crew continued with installing parts on the nose of DEIMOS. They are installing the nose cladding, but some of it will interfere with the installation of the electronics and will have to be deferred until later. The electronics team continues to connect systems. They expect to be ready to start testing some systems either tomorrow on early next week.

Dave's report from March 11-13: Jeff, Matt, and Jim installed the grating sliders. With the help of Pat Crivello, the kinematic mounts were modified. This was done by lowering the hard pads 3/8 of an inch and then lowering the rear kinematics on the instrument by a similar amount. The supports for the forward cladding were installed. Barry and Carol continued installing the electronics and are near the point where they will start to test things using the laptop computer. David Hilyard, Sandy, and Vernon continued with assembly of the camera. They centered element 3 and measured the tilt of camera lens group 4, which is acceptable. Body groups 1 through 4 have now been assembled, and they are preparing to test the camera. Jim Ward assembled the camera test stand, and the camera is installed in it. The first tests will be warm in the instrument lab. Following those tests the camera will be removed from the test stand, put in a box and athermalized in the dome. Cold testing in the dome will be done next week. Jack, Dick and Ken arrived in HI today.

Dave, March 7: The camera optics, assembly and test equipment arrived safely at the summit. David H, Vernon, and Sandy starting setting up in the instrument lab for assembly of the camera. They may actually start assembly tomorrow. Carol and Barry, with help from the mechanical crew, installed controllers 1 and 2 and started wiring them up. Carol also worked on installing the interconnect boxes for the stages. More parts of the nose structure were installed. DEIMOS was moved from its storage position to the observing position. The nose fits into the telescope without problem. Another big milestone passed. Pat Crivello's daughter wove a beautiful tea leaf lei for DEIMOS, which is greatly appreciated.

Dave, March 6: Keck connected DEIMOS to the glycol cooling system today, and the DEIMOS system is leak free. We can proceed installing the electronics panels. The structure that holds the TV camera was installed along with the front ring of the instrument. Barry and Carol continued unbundling and preparing cabling for the installation of the electronics panels. The camera arrived on the Big Island today. Everything seems to have gone well, and we expect to take the camera up to the observatory tomorrow. Arriving with the camera were Sandy, David Hilyard, Vernon, Jeff and Matt, making the team 9 people for the next couple of days. The snow is slowly going away, and, at least outside the dome, it was much warmer on the summit.

Dave, March 5: We installed the square tubes that hold the TV camera assembly and nose disk on the front of DEIMOS. Also installed was an additional part of the grating transport system. Barry and Carol worked on unbagging and uncoiling the cables in DEIMOS. Keck personnel are prepared to connect DEIMOS to the observatory glycol system tomorrow.

Dave, March 4: The hoist was installed in the instrument lab by Pat Crivello. Dick Kanto and Pat load tested it by lifting a 600 lb weight. Worked great! Carol and Barry unpacked the electronics and are ready to start working on cabling in the instrument tomorrow. Dick, with a little help from Pat and myself, installed two supports for the instrument nose structure, the support for the slit mask system, and one part of the grating slider rails. A good day!

Dave, March 1: The grating box, counter weight bar and weights, and the "grating system moving counter weight system" were installed on DEIMOS today. We also installed the DEIMOS mast which works with the cable boom that Keck installed on the platform. Jeff, Dick, and Matt completed their second week with this great effort. The Keck staff have been giving us wonderful support which is greatly appreciated.

Dave, Feb 28: DEIMOS rotates under manual power. We do not plan to install the motor until near the end of the DEIMOS assembly. We adjusted the kinematics, and DEIMOS now runs up on them rather than over the top. The kinematics are all at their nominal locations. We started to unpack the grating box and associated equipment and will be ready to lift them onto the platform tomorrow.

Dave, Feb 27: The DEIMOS storage position clamps were installed on the Nasymth platform. The shipping constraints that supported the drive disk during transportation were replaced with the spring-loaded cam followers. The cable wrap support was installed. We expect to be able to rotate DEIMOS tomorrow. DEIMOS was moved onto the deck and then on to its kinematics on the Nasymth platform. It cleared everything properly. The mounts still need to be aligned, but this was a great milestone to pass.

Dave: Feb 26 was a snow day. Snow extended down to 8,000 thousand feet of elevation below HP. They did not open the road to the summit until 4:30.

Dave, Feb 25: DEIMOS moved onto the telescope today. The telescope didn't seem to notice the extra 6 tons. There were no problems at all moving to the platform. DEIMOS is now being stored on the platform in NRT1. The forward restraints and the Destaco clamps were installed on the kinematics. The transition plates to the telescope that hold the Destaco clamps and forward restraints in the storage position, which were not fabricated previously, were completed. They are ready for installation onto the telescope tomorrow.

(Dave) Feb22: The air shipment and the marine container shipment arrived today, and all appear to be in good shape. The marine shipment is stored in the K2 dome, and the air shipment is in the Instrument lab. Jeff, Matt and Dick completed the installation of the kinematic mounts today, and we are ready to move DEIMOS onto the Nasymth Platform on Monday. We will not be able to test the kinematic mounts until Tuesday afternoon when NIRSPEC is moved out of the observing position.

(Dave) Feb 21: It was cold last night, and temperature control system in the dome kept it cold during the day (-3.2C). The four kinematic mount plates were installed in the platform and the back two kinematic mounts were installed on the plates. The back two kinematics could potentially interfere with NIRSPEC, which was moved into the observing position at noon today. The clearance was measured to be 3/8's of an inch. At the end of the day the control box and the cable reel for the tractor were installed on DEIMOS.

From Dave: It snowed last night. The road above HP was closed to the summit in the morning. While we were waiting for the road to open, Bob Moskitis convened a safety meeting to discuss DEIMOS with the entire summit crew. The road crew had the road open by about 10:30, and we arrived at the summit just after 11:00. Jeff, Matt T and Dick completed drilling and tapping the holes for the kinematics in the remaining part of the day.

It was foggy on the summit yesterday, so good thing we got the lift in early. This is my notes (Dave's) of the day's activities of Feb 19:
Arrived at the summit at 8:00
Installed the guides that keep DEIMOS on the rack
Installed the bracket that attaches the tractor to DEIMOS
Assembled the template that locates the DEIMOS kinematics on the Nas Platform
There is an interference with the forward gratings, particularly on the right side as you look into the telescope. The solution is to remove part of the grating. There is also a minor interference with the plate and the rail channel. The plates (front) will be slightly modified to accommodate this interference
At the end of the day one of the holes for one of the rear kinematics had drilled and tapped.

From Dave, 2/16:
On Thursday the crane was towed up the hill and set up outside the K2 dome. DEIMOS arrived at Kawaihae. It was met by the insurance surveyors (Mike Doyle and Randall Lund) and Matson's port manager, Russell Chin. Matson handled DEIMOS with great care, and none of the shock sensors had tripped. On Friday we did a dry run of the lift, and all went exactly as planned. Looks great for the lift on Sunday. On Sat, DEIMOS was towed up to the summit. Kona Trans did a great job, and got it there exactly when they said they would. The shock sensors were not tripped on this trip either. Tomorrow, weather permitting, we plan to lift into the dome.

From Sandy, 2/18:
It's nice to report that DEIMOS got into the dome safely yesterday and onto the Nasmyth platform. It was a bit touch-and-go with the weather. Dave Cowley said that winds were a bit higher than he would have liked, and bad weather is now closing in and will probably prevent a lift for at least another week. So we just made it, thanks to aggessive scheduling and prodding by Dave on the scene and to strong cooperation by CARA and Puna Crane. Congratulations to all involved.

The weather at Mauna Kea is perfect, according to Dave, so things are moving ahead of schedule. DEIMOS arrived on the big island safely at 4:30 with 2 surveyors on hand for inspection and observation (one hired by Matson, one hired by us). Dave's new plan is that DEIMOS gets picked up by the trucking company this morning, taken to their yard, and then trucked up to the summit Saturday morning. The lift into the dome will happen Sunday. The only snag right now is that Matson requires payment before they release items to the customer, so we are trying to accomodate this change in the schedule.

The main frame reached Honolulu successfully and was loaded up on the barge to Kawaihae. The marine surveyors in Oakland and in Hawaii say everything went smoothly, i.e., nothing was dropped and no shock indicators were set off. It should arrive in Kawaihae today. The closed container made it on the ship yesterday after a misunderstanding with the trucking company (6 AM, not PM!). The crane is going up to the summit today, and if weather holds, they'll do a test run tomorrow, 2/15. Since the weather is so good right now, the main frame will get trucked to the summit tomorrow, 2/15/02, and the actual lift into the dome might happen as soon as Saturday, 2/16/02.